OpenStudy (anonymous):

How do you find the maclaurin series for 10^x?

OpenStudy (zzr0ck3r):

\[\sum_{n=1}^{\infty}\frac{f^n(0)x^n}{n!}\]

OpenStudy (zzr0ck3r):

start taking the derivatives

OpenStudy (anonymous):

Like f'(x) = x10^(x-1)?

OpenStudy (zzr0ck3r):

f^1 = ln(10)*10^x f^2 = ln^2(10)10^x .... so the derivatives at x = 0 are ln(10) ln(10)^2 ln(10)^3 so the maclorian is \[\sum_{n=1}^{\infty}\frac{(ln(10)x)^n}{n!}\]

OpenStudy (anonymous):

No, that's not correct. The power rule applies for \[\frac{ d }{ dx }x^n\] but \[\frac{ d }{ dx } 10^x\] is something quite different! Learn how to differentiate exponential functions.

OpenStudy (zzr0ck3r):

no derivative of a^x = ln(a)*a^x

OpenStudy (zzr0ck3r):

don't worry everyone forgets how to differentiate that....common mistake to confuse it with a power rule.

OpenStudy (anonymous):

Thanks, so then, thats it?

OpenStudy (zzr0ck3r):

that was confusing;p

OpenStudy (zzr0ck3r):

d/dx e^x is the same ln(e)*e^x = 1*e^x = e^x

OpenStudy (zzr0ck3r):

so proving a^x with e^x does not help:P

OpenStudy (zzr0ck3r):

just memorize it

OpenStudy (anonymous):

I don't think I can forget it now, thanks!